Let B = {(1 2),(3 5)} and B'= {(1-1),(1-2)} be two bases of R2. Also, let...

  1. Let B = {(1 2),(3 5)} and B'= {(1-1),(1-2)} be two bases of R2. Also, let TB:  R2R2 be a linear operator on B described as TB(x,y) = (x+2y,x+y). Find TB'.
Answer #1

Here , B= {(1,2),0,03, B=f(),1),(1,-2)} - T: 18² 1R² is defined as (x,y) = (x+ 2y, sety To Rind mateix of T a.rt. B - I ((1,0I ((1,2)=(53) = 13 (1,-1)-8(1,-2) T(3,3)= (13,8)= 3461,-1) -21(1,-2). :. TBB= -8 -21 To find TBIR, I ((+1,-1)) =(4,0)=5(1,2)